# ScaleRAFT: Cross-Scale Recurrent All-Pairs Field Transforms for 3D Motion Estimation

### v1

@@ IV-A Datasets and Training Setting
Dataset. The training and testing of ScaleRAFT require the use of scene flow datasets, so we selected the currently commonly used KITTI[27] and Flyingthings3D[28] for Evaluating, and training based on Driving[28], Sintel[83], KITTI and Flyingthings3D. Because KITTI provides an official test ranking platform and contains the most advanced results, we mainly conduct evaluations on KITTI.
Training Setup. For the KITTI evaluation, we first conduct unified pre-training on the virtual datasets (Driving,Sintel, Flyingthings3D) and then fine-tune separately according to different evaluation items. Specifically, for the MID task, we use the same strategy as Yang[18] to split the dataset, selecting one out of every five images in KITTI for evaluation (K-40) and the remaining 160 images for fine-tuning training (K-160). For the scene flow task, we use all 200 images (K-200) for training and submit the results to the official evaluation platform for evaluation. During pre-training, the image size is 368x768, batch = 6, iteration = 200K, lr = 0.0005, and the number of motion foreground is 1. When fine-tuning, the image size is 960x320, batch = 6, lr = 0.000125, iterations = 80K, and the number of motion foreground is 1.
For the Flyingthings3D evaluation, we followed the training and evaluation process of RAFT3D[13], trained 200K generations on the training set (lr = 0.00025), and evaluated our method on the testing set. Only pixels with a depth of less than 35m and not occluded were included in the evaluation.
Figure 8: Motion-in-depth results for images “105” in the K-40 set. For each of them, Up: input images and the error map, where the whiter the pixels, the greater the error. Down: ground truth and visualization of motion-in-depth. Our method is much more accurate than other methods.
IV-B Motion-in-depth (MID)
MID is the core indicator of monocular 3D flow methods. In order to prove the superiority of ScaleRAFT in estimating MID tasks, we compared ScaleRAFT with the stereo-based traditional scene flow methods, state-of-art MID methods and scene flow methods on K-40. We use the same criteria as predecessors[18] to define the loss:
$Mid_{error}=||log(f_{3})-log(\bm{\tau}_{gt})||_{1}\cdot 10^{4}$ (26)
where $\bm{\tau}_{gt}$ is the ground truth of MID.
@@ IV-B Motion-in-depth (MID)
TABLE I: Motion-in-depth estimation on K40. The best among all are bolded, and the second best are underlined. Our method outperforms the monocular baselines by a large margin.
Method Input Training set $Mid_{error}(\downarrow)$ Time/s
@@
ScaleRAFT (ours) Mono K-160 42.08 0.2
RigidMask+ISF[15] RGB-D K-200 49.7 3.3
RAFT3D[13] RGB-D K-200 22.01 2
ScaleRAFT (ours) Mono K-200 29 0.2
We first compare with the stereo-based traditional approach: OSF and PRSM decompose the pixels into rigid blocks, and iteratively update the optical flow and depth field based on rigidity assumptions and hand-designed regularization terms. To get the MID result, we divide the depth of the second frame by the depth of the first frame. The experimental results show that our error is much smaller (115 v.s. 42.08) and in less time.
Then we compare with the state-of-the-art MID methods Optical expansion and Binary TTC. As shown in Fig. 8 and Tab. I, ScaleRAFT still has apparent advantages (73.5 v.s. 42.08). Unlike their direct regression of MID from optical flow, we embed the calculation of MID into optical flow matching. Convergence of information and end-to-end architecture make the final result more robust and accurate.
Finally, we compare ScaleRAFT with the state-of-the-art scene flow method. RigidMask and RAFT3D jointly optimize the calculated depth and optical flow field based on the rigidity assumption, and need to introduce additional semantic networks to calculate the rigidity mask. Because they used depth and optical flow networks pre-trained on KITTI, the training set is K-200. As shown in Tab. I, although they used additional input information, our method still has competitive accuracy (29 v.s. 22.01) and in less time, which reflects the effectiveness of our cross-scale matching strategy.
@@
TABLE II: State-of-the-art published methods on KITTI scene flow benchmark. RGB-D means depth and monocular information, and Mono means monocular information. D1, D2, Fl, and SF is the percentage of disparity, optical flow and scene flow outliers. -$bg$,-$fg$ and -$all$ represent the percentage of outliers averaged only over background regions, foreground regions and overall ground truth pixels. The best among all are bolded, and the second best are underlined. Our monocular method performs quite well, not only far ahead of similar monocular methods, but also outperforming RGBD-based methods in foreground scene flow estimation, which proves that the cross-scale matching can accurately estimate the 3D motion.
@@
CamLiFlow[14] RGB-D 1.48 3.46 1.81 - - 3.19 - - 4.05 - - 5.62
CamLiFlow-RBO[14] RGB-D 1.48 3.46 1.81 1.92 8.14 2.95 2.31 7.04 3.1 2.87 12.23 4.43
RigidMask+ISF[15] RGB-D 1.53 3.65 1.89 2.09 8.92 3.23 2.63 7.85 3.5 3.25 13.08 4.89
RAFT3D[13] RGB-D 1.48 3.46 1.81 2.51 9.46 3.67 3.39 8.79 4.29 4.27 13.27 5.77
ScaleRAFT3D (ours) RGB-D 1.48 3.46 1.81 1.93 8.16 2.97 2.37 9.26 3.51 2.91 13.87 4.74
ScaleRAFT-RBO (ours) RGB-D 1.48 3.46 1.81 1.93 7.72 2.89 2.27 5.63 2.83 2.86 10.91 4.20
@@
Binary TTC[19] Mono 1.48 3.46 1.81 3.84 9.39 4.76 5.84 8.67 6.31 7.45 13.74 8.5
RAFT[20] Mono - - - - - - 4.74 6.87 5.10 - - -
TPCV[86] Mono 1.48 3.46 1.81 2.29 7.63 3.18 4.53 5.52 4.69 5.34 10.6 6.21
Scale-flow[29] Mono 1.48 3.46 1.81 2.55 8.24 3.50 5.24 5.71 5.32 6.06 11.32 6.94
ScaleRAFT (ours) Mono 1.48 3.46 1.81 2.31 7.42 3.16 4.45 4.76 4.50 5.31 10.18 6.12
IV-C Time-to-collision (TTC)
Time-to-collision estimation is a crucial technique in path planning for autonomous robots[6, 7, 8, 87], which describes the collision time between moving objects and the observer plane. TTC can be estimated indirectly by MID:
where $T$ is the sampling interval between two consecutive frames, for KITTI $T=0.1s$, $Z$ and $Z^{\prime}$ are the object’s depth in frame1 and frame2, $\bm{\tau}$ is the motion-in-depth.
TABLE III: Percentage errors of time-to-collision estimation on K-40 validation set
Method Err-1s Err-2s Err-5s Time/s
Binary TTC (Binary mode)[19] 2.4 3.16 3.25 0.0192
Optical expansion[18] 1.86 1.87 2.45 0.2
OSF[50] 1.79 2.93 4.03 3000
ScaleRAFT (ours) 0.87 1.23 1.31 0.2
Referring to Binary TTC, we regard TTC as a binary classification task, judging whether the TTC of each pixel in the image is less than $\{1s,2s,5s\}$. We evaluate the percentage of outlier pixel values in prediction results and only consider pixels close to the observation plane.
Tab. III shows that ScaleRAFT achieves the lowest error in all time intervals, which proves our ScaleRAFT can estimate future collisions more accurately. As shown in Fig. 9, our method can effectively predict future collisions.
IV-D Scene Flow
@@
ScaleRAFT. We first compare ScaleRAFT with monocular methods, as shown in Tab. II, ScaleRAFT leads in all monocular methods with significant advantages, consistent with the previous superior performance in motion-in-depth tasks. It proves the effectiveness of our ScaleRAFT framework in the field of monocular 3D motion estimation. We also compare ScaleRAFT with the state-of-the-art RGBD-based methods RAFT3D, CamLiFlow, and RigidMask+ISF. As shown in Tab. II, RGBD-based methods tend to get smaller Fl-bg errors, which is thanks to the aid of depth-based rigid assumption to optimize the background optical flow. Nonetheless, our monocular approach still achieves the best results on SF-fg. Moreover, our method showed a significant improvement in optical flow prediction performance compared to the baseline model RAFT.
ScaleRAFT-RBO Although ScaleRAFT is not a framework specifically designed for RGBD, ScaleRAFT-RBO still has particular advantages over other state-of-the-art methods. As shown in Tab. II, the performance of ScaleRAFT-RBO exceeds most existing methods (4.43 v.s. 4.20). This excellent performance is mainly due to RBO compensating for the weakness of monocular methods in background estimation and CSCV alleviating the difficulty of foreground object matching.
ScaleRAFT3D. Compared with the baseline RAFT3D, ScaleRAFT3D has great advantages in all major indicators, especially the overall scene flow error SF-all has been dropped from 5.77 to 4.74, and has competitive accuracy in RGBD-based methods. We have noticed that the foreground part of ScaleRAFT3D has not improved. This is because RAFT3D is a system composed of multiple sub-modules, and the final result depends on the semantic-based rigid optimization pipeline. Strengthening a single 3D flow pipeline cannot comprehensively improve the performance of the baseline model.
TABLE IV: Performance comparison on the Flyingthings3D test set. We followed the setup of Flownet3D in which only non-occluded points with depths less than 35m are considered for evaluation. $\delta_{2D}<1px$ means the percentage of optical flow endpoint error less than 1 pixel, $\delta_{3D}<0.05m$ means the percentage of pixels with endpoint error less than 0.05m.
Method Input 2D Metrics 3D Metrics
$\delta_{2D}<1px(\uparrow)$ $EPE(\downarrow)$ $\delta_{3D}<0.05m(\uparrow)$ $\delta_{3D}<0.10m(\uparrow)$ $EPE(\downarrow)$
RAFT[20] Mono 77.2% 3.76 -- -- --
ScaleRAFT (ours) Mono 80.5% 3.41 31.7% 51.2% 0.206
FlowNet3D[53] LiDAR -- -- 25.4% 57.9% 0.169
FLOT[52] LiDAR -- -- 34.3% 64.3% 0.156
FlowNet3D++[89] RGB-D -- -- 30.3% 63.4% 0.137
RAFT3D[13] RGB-D 81.3% 2.43 82.1% 88.1% 0.064
ScaleRAFT3D (ours) RGB-D 85.4% 2.02 85.4% 90.8% 0.049
IV-E Flyingthings3D
To further verify the performance of our method, we have carried out more tests on the larger virtual dataset Flyingthings3D (Fly3D). We followed the configuration of RAFT3D[13] for training and evaluation. Specifically, trained 200k generations on the training set and evaluated our network with 2D and 3D endpoint errors (EPE). The 2D EPE is defined as the Euclidean distance between the real and predicted optical flow. 3D EPE is the European distance between the real and predicted scene flow.
In the testing phase, we used the split strategy proposed in FlowNet3D[53], containing 2007 test examples sampled from the Fly3D’s test set (Fly3D-2007). Only non-occluded pixels with depth $<$35m were included in the evaluation.
We first compared our ScaleRAFT3D with the baseline model RAFT3D. As shown in Tab. IV our expanded method has a leading advantage in all indicators, which proves that our scale matching strategy is superior to the regression strategy in RAFT3D.
We also explored the performance of our monocular method ScaleRAFT on the Fly3D. As shown in Tab. IV, ScaleRAFT has a leading performance in 2D indicators compared with RAFT, and has a certain competitiveness in 3D indicators compared with LiDAR and RGB-D methods, which proves that the ScaleRAFT framework can adapt to data containing challenging scenarios.
TABLE V: Ablation experiments. $Fl$-$epe$ is the end-to-end optical flow error, $M_{f}$ is the number of flight foregrounds.
Experiment Method $M_{f}$ $r_{s}$ $Fl$-$epe$ $Mid_{error}$
Scale Lookup Radius 0 0 3.52 160.63
0 $2/S$ 3.43 150.41
0 $1/S$ 3.47 147.92
Scale Feature W/o $F_{scale}$ 0 $1/S$ 3.49 226.78
W/o $L_{c}^{s}$ 0 $1/S$ 3.55 226.28
Hybrid Training Pipeline 1 $1/S$ 3.28 129.53
2 $1/S$ 3.22 132.38
3 $1/S$ 3.45 139.31
IV-F Ablation
In this section, we conduct ablation experiments to further confirm the ScaleRAFT structure’s rationality. We train ScaleRAFT 280K iterations on Driving and Flyingthings3D, with batch = 6, lr = 0.00025, and conduct testing on K200. Our algorithm’s default configuration is $r_{s}=1/S$, $M_{f}=1$.
Scale Lookup Radius. The scale lookup radius specifies the sampling range of lookup operator $L_{c}^{s}$, but does not change the scale correlation feature’s dimension. Generally, when $r_{s}$ is between $1/S$ and $2/S$, a slight change will not affect the final performance.
Scale Feature. We designed two ablation models to verify the effectiveness of scale lookup operator and scale correlation features.
In the “W/o $F_{scale}$” model, we removed the scale correlation feature $F_{scale}$ from the default model and only used optical flow correlation features $F_{of}$ to predict the 3D flow. As shown in Tab. V, the performance of both optical flow and MID has significantly decreased, and the accuracy of 3D metrics has decreased more (147.92$\rightarrow$226.78). It proves that scale correlation feature $F_{scale}$ is essential in depth motion estimation.
In the “W/o $L_{c}^{s}$” model, we do not use the scale lookup operator $L_{c}^{s}$, but directly use the $F_{multi}$ for motion estimation in the depth direction. The experimental results show that the error of the ablation model is still larger than the default model, even if more features are used.
Hybrid Training Pipeline. We also tested the effect of different numbers of foreground flying objects on the training results. As shown in Tab. V, when the current number of foregrounds is 1 and 2, performance significantly improved compared to the ground-truth training. When the current number of foregrounds reaches 3, the proportion of texture-matching learning exceeds regression learning, and the testing performance decreases.
TABLE VI: Breakdown of the inference latency. Results are computed on 1242x375 images with a RTX3090 GPU. The increased running time compared to the baseline model is highlighted in bold.
Models Component Individual Latency Total Latency
ScaleRAFT Feature extract 25ms (+20ms) 170.1ms
Construct CSCV 0.8ms (+65ms)
Context extract 4.3ms
GRU update 128ms (+33ms)
Last $f_{3}$ update 12ms (+12ms)
ScaleRAFT3D Feature extract 21.9ms (+17.6ms) 289.7ms
Construct CSCV 0.8ms (+17.6ms)
Resnet50 11ms
GRU update 216ms
$SE(3)$ update 40ms
V Latency and Limitations
Latency. In Tab. VI, we provide the inference breakdown time on KITTI. The timing results were measured on RTX3090. We calculated the 1242x375 size image 200 times consecutively, taking the average time as the final latency. As we can see, the introduction of CSCV adds a total time of 65ms to the RAFT baseline, with 20ms used to extract additional multi-scale features and 45ms used to calculate motion in the depth direction. For the 3D flow baseline RAFT3D, the introduction of CSCV only increased the total time by 6.1% (17.6ms), which is spent on extracting multi-scale features. Note that we do not use acceleration technology like TensorRT [^4] https://github.com/NVIDIA/TensorRT, so there is still room for optimization.
Limitations. Because ScaleRAFT needs to build a multi-scale cost volume, this increases memory usage from 3GB to 4.5GB compared to the RAFT baseline (only forward inference, with an input image size of 1242x375). Due to limited GPU memory, currently, we can only set the scale constant $S$ to 4. In addition, the depth clues provided by our ScaleRAFT are only accurate (error $<$ 5%) when the rotation speed is lower than 540°/s (default 30FPS video input), and the depth motion estimation of high-speed rotating objects may fail. Therefore, our method is more suitable for driving scenarios with less rotation.
VI Conclusion

### v2

@@ IV-A Datasets and Training Setting
We mainly compared the methods quantitatively on the KITTI[33] and Sintel[34] baselines, as they have publicly available online evaluation websites and have been widely applied in previous works.
Datasets Used: 200 image pairs from KITTI 2015 (K15)[33]. Commonly used pre-training datasets Flyingthings3D (T)[60], Driving (D)[60], Monkey (M)[60] and Sintel (S)[34] from movie.
Dataset Split: For the convenience of ablation experiments, we also divided the Sintel dataset into two subsets. One is the training set (S-18) containing 824 image pairs from 18 scenes, and the other is the testing set (S-5) containing 240 image pairs from 5 scenes. In addition, for the MID task, we use the same strategy as yang[20] to split the dataset, selecting one out of every five images in KITTI for evaluation (K-40) and the remaining 160 images for fine-tuning training (K-160).
Training Setting: Tab. I shows the specific training process, during which we also performed enhancement operations such as scaling, translation, and color domain transformation on the image pairs. Please refer to RAFT[22] for details.
IV-B Ablation
In this section, we discuss the effectiveness of the various components of the ScaleFlow++ method proposed in this paper. The complete ablation results are shown in Tab. II; we gradually stack the proposed modules based on the Base model.
TABLE II: Ablation Experiment. Resnet* represents the use of ImageNet weights for initialization. $Mid$ is the motion-in-depth error, and the calculation method is detailed in Eq. 24. EPE is the endpoint error of optical flow, and ALL is the outlier rate.
S-5-clean S-5-final K15
Ablation Iter EPE $\downarrow$ ALL $\downarrow$ EPE $\downarrow$ ALL $\downarrow$ EPE $\downarrow$ ALL $\downarrow$ $Mid\downarrow$
A Base 10k 2.70 11.43 3.23 15.01 4.89 18.95 167.86
B Base+Ht 10k 2.46 10.78 3.17 14.88 4.63 17.57 153.98
C Base+Ht+CSCV 10k 2.35 10.03 2.98 13.53 4.28 16.12 143.66
D Base+Ht+CSCV+Resnet 10k 2.35 10.00 3.17 14.07 4.38 16.47 137.56
E Base+Ht+CSCV+Resnet* 10k 2.11 8.42 2.81 11.69 4.96 17.13 175.84
F Base+Ht+CSCV+Resnet*+RNN 10k 2.06 8.11 2.63 11.76 3.51 11.85 127.55
G Base+Ht+CSCV+Resnet*+GIR 10k 1.87 7.59 2.53 10.51 3.60 11.91 115.03
H Base+Ht+CSCV+Resnet*+GIR+Init 10k 1.82 7.09 2.29 9.77 3.17 10.31 114.78
15k 1.72 6.66 2.23 9.24 3.23 10.20 112.73
I Base+Ht+CSCV+Resnet+GIR+Init 10k 2.04 8.32 2.77 11.66 3.15 10.83 95.52
15k 1.89 7.68 2.57 10.12 3.26 10.61 96.07
Base Model: We used the previous version Scale-flow as the ablation baseline and replaced the CSCV with the 4D Correlation Volume in RAFT. The rest remains unchanged, using the residual feature extractor and GRU iterative optimizer from RAFT.
Hybrid Training Pipeline (Ht): As shown in rows A and B of Tab. II, the mixed training pipeline Ht comprehensively improves the performance of all indicators. This proves that our proposed hybrid training pipeline can effectively enhance the scene flow performance of the network without increasing training costs.
Cross-Scale Matching (CSCV): As shown in rows B and C of Table II, the model’s optical flow and scale flow performance have been significantly improved. Especially in the driving scenario KITTI with significant changes in scale, the error of optical flow decreased from 4.63 to 4.28, which proves the effectiveness of our cross-scale matching strategy.
ResNet and Weight Pre-training: In rows D and E of Table II, we replaced the original residual convolution with standard Resnet and used pre-trained weights from ImageNet. The experimental results show that the performance of Resnet is close to the accuracy of the RAFT original feature extractor, but after using pre-trained weights, the performance of the Sintel dataset is significantly improved, while the KITTI dataset is significantly reduced. This is easy to understand because there are almost no driving scenes in ImageNet, and its data domain is closer to Sintel. We also verified the relationship between training iteration and pre-training weights. As shown in H and I, the method using pre-training weights still maintained significant advantages when more iterations were being trained.
The above experiment proves that appropriate pre-training weights can effectively improve model performance in similar data domains, and this advantage will continue to be maintained with increasing iteration times. Therefore, we recommend using ImageNet weights only on the Sintel dataset in subsequent comparisons.
Figure 8: Initialization Module. We selected the first 100 sequences of K15 to calculate the average optical flow endpoint error (EPE) for each iteration. The blue and red colors in the figure correspond to H and G in Table II, respectively. The difference is that the red has an additional GIR iteration to ensure fair comparison. When the optical flow field is all 0, the initial EPE error is 36.72, and the initialization module reduces the initial error to 10.96, reducing the subsequent optimizer’s training difficulty and achieving higher accuracy with fewer iterations.
TABLE III: Runtime Analysis of Different Optimizers. Perform 100 consecutive inferences on a $375\times 1242$ image and take the average. The GPU used is RTX4090.
Optimizer Single Iteration Time Refine Iterations Total Time
GIR (ours) 3.26ms 6 19.56ms
RNN 1.92ms 12 23.04ms
GRU 1.62ms 12 19.44ms
Iterative Optimizer: We compared the performance between the most commonly used GRU optimizer, RNN optimizer in SEA-RAFT[58], and the GIR optimizer proposed in this paper. As shown in rows E, F, and G in Table II, all optimizers have undergone 6 iterations, and GIR has significant advantages in Sintel scenarios and $Mid$ metrics. Moreover, we also reported the running time of each optimizer in the complete method in Table III. The optimizer proposed in this paper did not significantly increase computational overhead under standard iteration numbers.
Initialize Displacement Field: As shown in row H of Table II, the initialization module significantly improved the performance of optical flow and slightly enhanced $Mid$ performance, proving the effectiveness of the initialization module.
We also demonstrated a more detailed working process of the initialization module. As shown in Fig. 8, the initialization module reduced the initial optical flow EPE from 36.72 to 10.96, alleviating the task alienation of the optimizer at different iteration stages, reducing the overall number of iterations, and improving accuracy.
@@ IV-B Ablation
TABLE IV: Motion-in-depth estimation on K-40. The best among all are bolded, and the second best are underlined. Our method outperforms the monocular baselines by a large margin.
Method Input Training set $Mid\downarrow$ Time/s
@@
TPCV [62] Mono K-160 42.84 0.2
ScaleFlow++(ours) Mono K-160 38.44 0.2
Figure 9: Motion-in-depth results for images “105” and “195” in the K-40 set. For each of them, Up: input images and the error map, where the whiter the pixels, the greater the error. Down: ground truth and visualization of motion-in-depth. Our method is much more accurate than other methods.
@@
IV-C Motion-in-depth (MID)
MID is the core indicator of monocular 3D flow methods. In order to prove the superiority of ScaleRAFT in estimating MID tasks, we compared ScaleFlow++ with the stereo-based traditional scene flow methods, state-of-the-art MID methods and scene flow methods on K-40. We use the same criteria as predecessors[20] to define the loss:
$Mid=||log(f_{3})-log(\bm{\tau}_{gt})||_{1}\cdot 10^{4}$ (24)
where $\bm{\tau}_{gt}$ is the ground truth of MID.
We first compare with the stereo-based traditional approach: OSF and PRSM decompose the pixels into rigid blocks, and iteratively update the optical flow and depth field based on rigidity assumptions and hand-designed regularization terms.To get the MID result, we divide the depth of the second frame by the depth of the first frame. The experimental results show that our error is much smaller (115 v.s. 38.44) and in less time.
Then we compare with the state-of-the-art MID methods Optical expansion, TPCV and Binary TTC. As shown in Table IV, ScaleFlow++ still has apparent advantages (42.84 v.s. 38.44).
Application example: We present an application example of MID in downstream 3D tasks: time-to-collision (TTC).
Time-to-collision estimation is a crucial technique in path planning for autonomous robots[64, 2, 10, 4], which describes the collision time between moving objects and the observer plane. TTC can be estimated indirectly by MID $\bm{\tau}$:
where $T$ is the sampling interval between two consecutive frames, for KITTI $T=0.1s$, $Z$ and $Z^{\prime}$ are the object’s depth in frame1 and frame2.
We demonstrate the TTC performance of ScaleFlow++ in Fig. 10. Our method can easily distinguish the collision sequence and approximate collision time of different moving objects.
IV-D Optical Flow
Results are shown in Tab. VI. Compared to the previous version of Scale-flow, ScaleFlow++ has improved the overall performance by at least 18% and 19% on Sintel and KITTI, respectively, demonstrating the effectiveness of the improvements proposed in this paper.
Moreover, compared with existing state-of-the-art optical flow methods, ScaleFlow++ is also very competitive. On the Sintel test, it leads most RAFT-based methods and achieves even better performance on the KITTI test.
TABLE V: State-of-the-art published methods on KITTI scene flow benchmark. RGB-D means depth and monocular information, and Mono means monocular information. D1, D2, Fl, and SF is the percentage of disparity, optical flow and scene flow outliers. -bg,-fg and -all represent the percentage of outliers averaged only over background regions, foreground regions and overall ground truth pixels. The best among the same group are bolded, and the second best are underlined. Our monocular method performs quite well, not only far ahead of similar monocular methods, but also outperforming RGBD-based methods in foreground scene flow estimation, which proves that the cross-scale matching can accurately estimate the 3D motion.
@@ IV-D Optical Flow
CamLiFlow[12] RGBD 1.48 3.46 1.81 - - 3.19 - - 4.05 - - 5.62
CamLiFlow-RBO[12] 1.48 3.46 1.81 1.92 8.14 2.95 2.31 7.04 3.1 2.87 12.23 4.43
RigidMask+ISF[49] 1.53 3.65 1.89 2.09 8.92 3.23 2.63 7.85 3.5 3.25 13.08 4.89
RAFT3D[39] 1.48 3.46 1.81 2.51 9.46 3.67 3.39 8.79 4.29 4.27 13.27 5.77
@@
Binary TTC[21] 1.48 3.46 1.81 3.84 9.39 4.76 5.84 8.67 6.31 7.45 13.74 8.5
TPCV[62] 1.48 3.46 1.81 2.29 7.63 3.18 4.53 5.52 4.69 5.34 10.60 6.21
Scale-flow[35] 1.48 3.46 1.81 2.55 8.24 3.50 5.24 5.71 5.32 6.06 11.32 6.94
ScaleFlow++ (ours) 1.48 3.46 1.81 2.13 8.02 3.11 3.94 5.59 4.21 4.81 10.69 5.79
TABLE VI: Compare optical flow performance with other state-of-the-art methods. The best among all are bolded, and the second best are underlined. * denotes the use of TartanAir[65] as additional training data.
Sintel-test K15-test
[…]
RAFT SEA-RAFT(M)* 1.44 2.86 4.64 4.47 5.49
SEA-RAFT(L)* 1.31 2.60 4.30 4.08 5.37
RAFT[22] 1.61 2.86 5.10 4.74 6.87
CRAFT[66] 1.45 2.42 4.79 4.58 5.85
MS-RAFT[67] 1.37 2.67 4.58 4.88 6.38
GMA[13] 1.39 2.47 5.15 - -
Scale-flow[35] 1.88 3.16 5.32 5.24 5.71
ScaleFlow++(ours) 1.33 2.77 4.21 3.94 5.59
Transformer GMFlowNet[68] 1.39 2.65 4.79 4.39 6.84
GMFlow[69] 1.74 2.90 9.32 9.67 7.57
FlowFormer[70] 1.20 2.12 4.68 4.37 6.18
GMFlow++[71] 1.03 2.37 4.49 4.27 5.60
Figure 11: Visualization results on the KITTI test. In each group, from top to bottom, there are two consecutive frames of images, the visualized optical flow, and the visualized optical flow error map. The redder the color of the error map, the greater the error. Observing the red box area in the image, our improved method better optimizes the background over a large area.
IV-E Scene Flow
@@
We first compare ScaleFlow++ with monocular methods, as shown in Table V, ScaleFlow++ leads in all monocular methods with significant advantages, consistent with the previous superior performance in motion-in-depth tasks. It proves the effectiveness of our ScaleFlow++ framework in the field of monocular 3D motion estimation.
@@
We also compare ScaleFlow++ with the state-of-the-art RGBD-based methods RAFT3D, CamLiFlow, and RigidMask+ISF. As shown in Tab. V, RGBD-based methods tend to get smaller Fl-bg errors, which is thanks to the aid of depth-based rigid assumption to optimize the background optical flow. Nonetheless, our monocular method still achieved competitive results in the core indicator SF-all and significant advantages in foreground estimation (SF-fg) compared to the RGBD-based method (10.69 v.s. 12.23).
Figure 12: Self-supervised generalization. ScaleFlow++ has demonstrated stunning results in natural scenes that have never been seen before, demonstrating that our method is robust and easy to generalize.
IV-F Motion-in-depth on KITTI Test
Since the KITTI testing platform does not directly indicate motion-in-depth, we indirectly measure the accuracy of motion-in-depth based on the depth error growth (DEG) rate.
$\textrm{DEG}=\frac{\textrm{D2}-\textrm{D1}}{\textrm{D1}}$ (26)
where D1 and D2 are the depth outlier rates in the first and second frames, respectively. DEG describes how much error is added to the depth corresponding to the second frame based on the depth estimation of the first frame.
TABLE VII: Comparison of Growth Rates of Deep Motion Error on KITTI Test.
Method Input DEG-$\downarrow$ DEG-$\downarrow$ DEG-$\downarrow$
CamLiFlow-RBO RGBD 0.30 1.35 0.63
RigidMask+ISF 0.37 1.59 0.71
RAFT3D 0.70 1.73 1.03
Optical expansion Mono 1.29 1.47 1.35
Binary TTC 1.59 1.71 1.63
TPCV 0.55 1.21 0.76
Scale-flow 0.72 1.38 0.93
ScaleFlow++ 0.44 1.32 0.72
As shown in Tab. VII, our method outperforms all monocular methods in DEG-all and is highly competitive even compared to RGBD methods. In addition, the performance of ScaleFlow++ is particularly impressive in the evaluation of the background area, mainly due to the better global view provided by the GIR module to optimize the background.
IV-G Generalization in Real World
To assist readers in better applying ScaleFlow++, we trained generalization weights on 60000 real-world image pairs using a 3D flow self-supervised generalization method, ADFactory[59].
As shown in Fig. 12, whether it is optical flow or MID, the ScaleFlow++ trained by self-supervision has shown excellent generalization results in unseen scenarios. We hope that our work can help methods in fields such as human pose recognition, action prediction, and autonomous driving better perceive the three-dimensional world.
V Conclusion